ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

1. Logic

1991 MS Classification: 03 G 25 - 08 B 99 Niederkorn, Philippe, Université de Liège Natural dualities for finitely generated varieties of MV-algebras A (generalized) natural duality is obtained for each finitely generated variety of MV-algebras. An axiomatization of the dual classes is given, as well as some applications.